An ice-making machine operates in a Carnot cycle. It takes heat from water at 0.0°C and rejects heat to

a room at 24.0°C. Suppose that 85.0 kg of water at 0.0_C are converted to ice at 0.0°C. (a) How much

heat is discharged into the room? (b) How much energy must be supplied to the device?


1
Expert's answer
2021-05-17T18:57:27-0400

answer:-

a)

QC = heat from ice , QH heat discards to the room

Lf= heat fusion ="334\times 10^3 J/kg"


"Q_c=mL_f=85kg\times334\times 10^3 J/kg=1.84\times10^4kJ"


now

"\frac{Q_C}{Q_H}=-\frac{T_C}{T_H}\\
{Q_C}=-{Q_H}\frac{T_C}{T_H}=-(2.84\times10^4kJ)(\frac{297K}{273K})\\
Q_c=\boxed{-3.09\times10^4kJ}"

the negative sign is due to discarding the heat to the hot reservoir


b)

"|Q_H|=Q_c+|W|
\\ |W|=Q_C-|Q_H|=2.84\times10^4-3.09\times10^3kJ\\
W=-2.50\times10^3kJ"